# Dataset Distillation for Offline Reinforcement Learning

### v2

@@ 2.1 Offline reinforcement learning problem setting
In our reinforcement learning setting, the environment is modelled as a Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},T,R,s_{0},s_{-1}\rangle$ with state space $\mathcal{S}$, action space $\mathcal{A}$, $T(s_{t+1}|s_{t},a)$ is the probabilistic transition function, $R(s_{t},a)\in\mathbb{R}$ is the immediate transition reward, $s_{0}$ is the start state, and $s_{-1}$ is the end state since we are in an episodic setting. A policy function $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is a mapping from a state to a probability distribution over the action space. We will assume that both the $\mathcal{S}$ and $\mathcal{A}$ are discrete in our setting. When a policy can be parameterized by some parameter $\theta$, we denote the policy as $\pi_{\theta}$. Since we are using deep reinforcement learning, $\pi_{\theta}$ will be a neural network with weights $\theta$.
@@ 2.2 Behavioral cloning
$\mathcal{L}_{\textbf{BC}}(\theta|\mathcal{D})=\sum_{(s_{t}^{i},a_{t}^{i},...)\in\mathcal{D}}w(s_{t}^{i},a_{t}^{i},...)||\pi_{\theta}(s_{t}^{i})-a_{t}^{i}||$ (1)
where ||||| | is some norm implying some notion of distance and $w$ is some weighting on the data-points in $\mathcal{D}$. Usually we take the uniform weighting $w(*)=1$. Since we usually use stochastic gradient descent (SGD) to minimize the loss, in this case we can simply use $p(*)=\frac{w(*)}{\sum_{(s_{t}^{i},a_{t}^{i},...)\in\mathcal{D}}w(s_{t}^{i},a_{t}^{i},...)}$ as the probability of selecting the sample.
Since $\pi_{\beta}$ might not be optimal, we can attempt to train a better policy $\pi_{\theta}$ by filtering out observations in $\mathcal{D}$ that lead to poor outcomes. In other words, we let $w(s_{t}^{i},a_{t}^{i},...,G_{0}^{i})=\mathbb{I}_{G_{0}^{i}\geq b}$ where $b$ is some threshold on good vs bad outcomes. We call the method of choosing $b$ such that we only end up with xof the initial dataset, and training a policy $\pi_{\theta}$ using BC on it as BCx%, or percentile behavior cloning. In other words, the goal of percentile behavior cloning is to filter out a better training dataset.
@@ 3.1 Environment
In Procgen environments, the state space, consists of 64x64 pixel images (RGB array with three channels and values 00 to $255$). The action space is discrete in nature and usually includes movements (up, down, left, right) and interactions like collecting items or opening doors. For our experiments, we consider three procedurally generated games: Bigfish, Starpilot, and Jumper. In Starpilot for example, the player must navigate a space ship to avoid being hit by bullets and shoot down enemies in an arcade game fashion. Enemies and obstacles are procedurally generated, so each ‘map’ is different. The key characteristic of the Procgen benchmark is that, given a different seed, the player encounters a different ‘map’ though the game rules are unchanged. A key challenge for AI agents lies in how well they can adapt to seeds that they have not seen before.

### v3

@@ 2.1 Offline reinforcement learning problem setting
In our reinforcement learning setting, the environment is modelled as a Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},T,R,s_{0},s_{-1}\rangle$ with state space $\mathcal{S}$, action space $\mathcal{A}$, $T(s_{t+1}|s_{t},a)$ is the probabilistic transition function, $R(s_{t},a)\in\mathbb{R}$ is the immediate transition reward, $s_{0}$ is the start state, and $s_{-1}$ is the end state since we are in an episodic setting. A policy function $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is a mapping from a state to a probability distribution over the action space. We will assume that both the $\mathcal{S}$ and $\mathcal{A}$ are discrete in our setting. When a policy can be parameterized by some parameter $\theta$, we denote the policy as $\pi_{\theta}$. Since we are using deep reinforcement learning, $\pi_{\theta}$ will be a neural network with weights $\theta$.
@@ 2.2 Behavioral cloning
[…]
where $||$ is some norm implying some notion of distance and $w$ is some weighting on the data-points in $\mathcal{D}$. Usually we take the uniform weighting w()=1w(*)=1. Since we usually use stochastic gradient descent (SGD) to minimize the loss, in this case we can simply use $p(*)=\frac{w(*)}{\sum_{(s_{t}^{i},a_{t}^{i},...)\in\mathcal{D}}w(s_{t}^{i},a_{t}^{i},...)}$ as the probability of selecting the sample.
Since $\pi_{\beta}$ might not be optimal, we can attempt to train a better policy $\pi_{\theta}$ by filtering out observations in $\mathcal{D}$ that lead to poor outcomes. In other words, we let $w(s_{t}^{i},a_{t}^{i},...,G_{0}^{i})=\mathbb{I}_{G_{0}^{i}\geq b}$ where $b$ is some threshold on good vs bad outcomes. We call the method of choosing $b$ such that we only end up with x% of the initial dataset, and training a policy $\pi_{\theta}$ using BC on it as BCx%, or percentile behavior cloning. In other words, the goal of percentile behavior cloning is to filter out a better training dataset.
@@ 3.1 Environment
In Procgen environments, the state space, consists of 64x64 pixel images (RGB array with three channels and values 0 to $255$). The action space is discrete in nature and usually includes movements (up, down, left, right) and interactions like collecting items or opening doors. For our experiments, we consider three procedurally generated games: Bigfish, Starpilot, and Jumper. In Starpilot for example, the player must navigate a space ship to avoid being hit by bullets and shoot down enemies in an arcade game fashion. Enemies and obstacles are procedurally generated, so each ‘map’ is different. The key characteristic of the Procgen benchmark is that, given a different seed, the player encounters a different ‘map’ though the game rules are unchanged. A key challenge for AI agents lies in how well they can adapt to seeds that they have not seen before.